OpenStudy (anonymous):

Compute the sum: S=1!.1+2!.2+3!.3+....+2011!.2011.

OpenStudy (zarkon):

(2012)!-1

OpenStudy (zarkon):

\[S=1!\cdot1+2!\cdot2+3!\cdot3+....+n!\cdot n=(n+1)!-1\]

OpenStudy (asnaseer):

yes - if we set the nth term as:\[n*n!\]then the term before this is:\[n!-(n-1)!\]so I believe a lot of the terms cancel out leaving you with the @Zarkons answer

OpenStudy (asnaseer):

which comes from expansion of:\[(n-1)!*(n-1)\]

OpenStudy (waheguru):

http://openstudy.com/#/updates/4ececc98e4b04e045af1b27a esay question

OpenStudy (zarkon):

it is a simple proof by induction

OpenStudy (waheguru):

U GUYS PLEASE HELP ME

OpenStudy (asnaseer):

I thought for induction you had to be given an equation upfront and then prove its true?

OpenStudy (zarkon):

I only know it of hand because I just assigned it this semester to my discrete math class

OpenStudy (zarkon):

and I did give you an equation ;)

OpenStudy (asnaseer):

where? I am not in any maths class - I just do maths as a hobby

OpenStudy (zarkon):

this is one of those problems where I would have looked at smaller numbers to see if i could find a pattern

OpenStudy (asnaseer):

or are you saying that the original question was set as prove ...=(n+1)!-1

OpenStudy (zarkon):

yes

OpenStudy (asnaseer):

ah - ok - then induction would work - thanks

OpenStudy (asnaseer):

but looking at the terms, I believe they can be written out in reverse order (i.e. starting from the nth term) as: (n*n!) + (n!-(n-1)!) + ((n-1)!-(n-2)!) + ...

OpenStudy (asnaseer):

so all the intermediate factorials cancel out

OpenStudy (zarkon):

yes

OpenStudy (asnaseer):

leaving: n*n! + n! - 1 = n!(n+1) - 1 = (n+1)! - 1

OpenStudy (asnaseer):

ok - so pattern recognition works too :-)

OpenStudy (zarkon):

almost always more than one way to do things

OpenStudy (asnaseer):

does pattern recognition count as a proper proof in maths?

OpenStudy (zarkon):

usually....but I asked for a proof by induction in my particular class

OpenStudy (asnaseer):

I was just curious as I am interested in mathematics and was wondering how "strict" proofs have to be before they are considered "worthy"

OpenStudy (zarkon):

doing it your way is find....I covered induction and I wanted to make sure that they new how to do it using indiction

OpenStudy (zarkon):

*fine

OpenStudy (asnaseer):

ok - thanks again for taking the time to respond.

OpenStudy (zarkon):

no problem :)

OpenStudy (anonymous):

You could make the sum into a telescoping sum too:\[\sum_{k=1}^{2011}k\cdot k!=\sum_{k=1}^{2011}((k+1)-1)\cdot k!=\sum_{k=1}^{2011}(k+1)!-k!=2012!-1!\]

OpenStudy (zarkon):

yes...that is what asnaseer did.

OpenStudy (anonymous):

ah my bad, i didnt take the time to read it all >.<

OpenStudy (asnaseer):

although I ever realised its called a telescoping sum :-)

OpenStudy (asnaseer):

**never

OpenStudy (asnaseer):

no probs @joemath - your posting has just incremented my maths vocabulary :-)

myininaya (myininaya):

zarkon you made a correction to one word in that sentence you should had made two corrections new should have been knew and you found the one with find which was suppose to be fine i just thought i would correct you on something since it makes me feel good lol